OpenStudy (anonymous):

What is the solution to the equation log3 x – log 3 2 = 2 ?

OpenStudy (anonymous):

\(\large log_3 x – log_3 2 = 2 \rightarrow log_3(\frac{x}{2})=2 \Leftrightarrow 3^2=\frac{x}{2}\) can you solve for x now?